### v2

# On Glaisher’s Partition Theorem

Abstract.
Glaisher’s theorem states that the number of partitions of $n$ into parts which repeat at most m1m-1 times is equal to the number of partitions of $n$ into parts which are not divisible by $m$. The $m=2$ case is Euler’s famous partition theorem. Recently, Andrews, Kumar, and Yee gave two new partition functions $C(n)$ and $D(n)$ related to Euler’s theorem. Lin and Zhang extended their result to Glaisher’s theorem by generalizing $C(n)$. We generalize $D(n)$, prove an analogous partition identity for the $m=3$ case, and show that the general case is an example of an almost partition identity. We also provide a new series equal to Glaisher’s product both in the finite and infinite cases.
Key words and phrases:
Partitions, Partition Identities, Euler’s Theorem, Glaisher’s Theorem, $q$-series, Almost Partition Identities.
2020 Mathematics Subject Classification:
05A17, 05A19, 11P81, 11P84.
1. Introduction
Euler’s celebrated partition theorem [1] can be stated as
(1.1) $A(n)=B(n)$
where $A(n)$ is the number of partitions of $n$ into distinct parts and $B(n)$ is the number of partitions of $n$ into odd parts.
In a recent work, Andrews, Kumar, and Yee [4] have provided two more elegant companion partition functions to $A(n)$ and $B(n)$. In particular, they have obtained the following partition identity.
Theorem $1.$
([4, Theorem $1.1$]) For any positive integer $n$, we have
(1.2) $A(n)=B(n)=C(n+1)=\dfrac{1}{2}D(n+1),$
where $C(n)$ is the number of partitions of $n$ with largest part even and parts not exceeding half of the largest part are distinct and $D(n)$ is the number of partitions of $n$ into non-negative parts wherein the smallest part appear exactly twice and no other parts are repeated.
Let $m\geq 2$ be an integer. Define $A_{m}(n)$ to be the number of partitions of n into parts repeating less than $m$ times and $B_{m}(n)$ to be the number of partitions of $n$ into parts not divisible by $m$. In 1883, Glaisher [5] obtained the following partition theorem which is a natural extension of Euler’s theorem (1.1).
Theorem 1.2.
([5]) For any non-negative integer $n$, we have
(1.3) $A_{m}(n)=B_{m}(n).$
For any integer $0\leq j\leq m-1$, define $B^{(j)}_{m}(n)$ to be the number of partitions of $n$ into parts not divisible by $m$ and largest part congruent to $j$ modulo $m$. Clearly,
(1.4) $B_{m}(n)=\sum\limits_{j=1}^{m-1}B^{(j)}_{m}(n).$
In a recent work, Lin and Zhang [6] generalized Andrews, Kumar, and Yee’s $C(n)$ function to the case of $m$-regular partitions (partitions whose parts are not divisible by $m$). It is noteworthy to point out that the cardinality of the set $B^{\prime}_{m}(n)$ in their paper is the same as $B^{(m-1)}_{m}(n)$.
Define $C_{m}(n)$ to be the number of partitions of $n$ in which the largest part is divisible by $m$ (say mjmj) and all parts less than or equal to $j$ repeat less than $m$ times. Also, define $C_{m}(0):=1$.
Theorem 1.3.
([6, Theorem $1.5$]) For any positive integer $n$, we have
(1.5) $B^{(m-1)}_{m}(n)=C_{m}(n+1).$
Remark 1.
Lin and Zhang [6] provided both $q$-series (through generating functions) and bijective proofs of Theorem 1.3.
Let us now recall some notations from the theory of $q$-series which can be found in [1]. For any complex variable AA and non-negative integer NN, the conventional $q$-Pochhammer symbol is defined as
$(A)_{N}=(A;q)_{N}$ $:=\Bigg\{\begin{array}[]{lr}1\qquad\qquad\quad\,\,\,\text{when }N=0,\\ \prod\limits_{i=0}^{N-1}(1-Aq^{i})\,\,\,\text{otherwise, and}\end{array}$
$(A)_{\infty}=(A;q)_{\infty}$ $:=\lim\limits_{N\rightarrow\infty}(A)_{N}\,\,\text{for}\,\,|q|<1.$
Next, for non-negative integers $a$, $b$, we define the $q$-binomial coefficient as
$\left[\begin{matrix}a+b\\ b\end{matrix}\right]_{q}:=\Bigg\{\begin{array}[]{lr}\dfrac{(q)_{a+b}}{(q)_{a}(q)_{b}}\quad\text{for }a,b\geq 0,\\ 0\qquad\qquad\text{otherwise}.\end{array}$
Also, let $\zeta_{n}=e^{\frac{2\pi i}{n}}$ denote the primitive $n^{\text{th}}$ root of unity and $\omega=\zeta_{3}$ denote the primitive cubic root of unity here and throughout the remainder of the paper.
Now, define $D_{m}(n)$ to be the number of partitions of $n$ into non-negative parts where the smallest part occurs exactly $m$ times and all other parts repeat less than $m$ times.
Theorem 1.4.
For any non-negative integer $n$, we have
(1.6) $C_{m}(n)=\dfrac{D_{m}(n)+E_{m}(n)}{m},$
where $E_{m}(n)$ is such that the generating function of $E_{m}(n)$ is
(1.7) $\varepsilon_{m}(q):=\sum\limits_{n=0}^{\infty}E_{m}(n)q^{n}=\sum\limits_{n=0}^{\infty}q^{mn}(q^{n+1};q)_{\infty}\sum\limits_{j=1}^{m-1}(\zeta^{j}_{m}q^{n+1};q)_{\infty}.$
The $m=3$ case of (1.7) corresponds to the following identity
Theorem 1.5.
(1.8) $\varepsilon_{3}(q):=\sum\limits_{n=0}^{\infty}E_{3}(n)q^{n}=2-q-2q^{2}+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n-1)q^{\binom{n+1}{2}+1},$
where $\chi_{3}(n)$ is defined as
$\chi_{3}(n):=\omega^{n}+\omega^{-n}=\Bigg\{\begin{array}[]{lr}2\qquad\,\,\,\text{if }3\mid n,\\ -1\quad\,\,\,\,\text{if }3\nmid n.\end{array}$
The complexity of the generating functions involved seemed to mitigate against an analog of the $C(n)=\frac{1}{2}D(n)$ portion of Theorem 1.1; however, surprisingly, a similar result exists for $m=3$ which clearly follows from (1.6) and (1.8).
Theorem 1.6.
Let $n$ be any positive integer that is not equal to a triangular number plus 11, then
(1.9) $C_{3}(n)=\frac{1}{3}D_{3}(n).$
Even more strikingly, we obtain another example of an almost partition identity (they are true for almost all $n$ but fail for an infinity of values; see [3]) for the general case which is as follows.
Theorem 1.7.
The identity
(1.10) $C_{m}(n)=\frac{1}{m}D_{m}(n)$
is true for almost all $n$. Moreover, if $N(x)$ is the number of times that (1.10) is true for $n<x$, then
$\lim\limits_{x\rightarrow\infty}\frac{N(x)}{x}=1.$
Combining Theorems 1.2, 1.3, and 1.4, we then have the following partition identity which is a natural extension of Theorem $1.$
Theorem 1.8.
For any positive integer $n$, we have
(1.11) $A_{m}(n)=B_{m}(n)=\sum\limits_{k=1}^{m-2}B^{(k)}_{m}(n)+C_{m}(n+1)=\sum\limits_{k=1}^{m-2}B^{(k)}_{m}(n)+\dfrac{D_{m}(n+1)+E_{m}(n+1)}{m}.$
In view of the generating function of $B_{2}(n)$ in [4, $(2.1)$], we also obtain the following finite $q$-series identity.
Theorem 1.9.
For any positive integer NN, we have
(1.12) $1+\sum\limits_{n=1}^{N}\sum\limits_{j=1}^{m-1}\dfrac{q^{mn-j}}{\left(\prod\limits_{r=1}^{m-j}(q^{r};q^{m})_{n}\right)\cdot\left(\prod\limits_{r=m-j+1}^{m-1}(q^{r};q^{m})_{n-1}\right)}=\dfrac{(q^{m};q^{m})_{N}}{(q;q)_{mN}}.$
Letting $N\rightarrow\infty$ in (1.12) gives us the following $q$-series identity.
Corollary 1.10.
(1.13) $1+\sum\limits_{n=1}^{\infty}\sum\limits_{j=1}^{m-1}\dfrac{q^{mn-j}}{\left(\prod\limits_{r=1}^{m-j}(q^{r};q^{m})_{n}\right)\cdot\left(\prod\limits_{r=m-j+1}^{m-1}(q^{r};q^{m})_{n-1}\right)}=\dfrac{(q^{m};q^{m})_{\infty}}{(q;q)_{\infty}}.$
Remark 2.
The $m=2$ case of (1.13) above was recently proved by Andrews and Bachraoui in [2, Lemma 11, (1.10)]. Clearly, the series on the left-hand side of (1.13) is the generating function of $\sum\limits_{j=1}^{m-1}B^{(j)}_{m}(n)$ and the product on the right-hand side of (1.13) is the generating function of $B_{m}(n)$. The same observation (restricting all parts to be at most mNmN) proves Theorem 1.9.
The rest of the paper is organized as follows. In Sections 2, 3, and 4, we provide $q$-series (generating function) proofs of Theorem 1.4, 1.5, and 1.7 respectively.
2. Proof of Theorem 1.4
From the definition of $C_{m}(n)$, we then have
$\sum\limits_{n=0}^{\infty}C_{m}(n)q^{n}$ $=\sum\limits_{n=0}^{\infty}\dfrac{\prod\limits_{j=1}^{n}(1+q^{j}+\ldots+q^{(m-1)j})q^{mn}}{(q^{n+1};q)_{(m-1)n}}$
$=\sum\limits_{n=0}^{\infty}\dfrac{(q^{m};q^{m})_{n}q^{mn}}{(q;q)_{mn}}$
(2.1) $=(q^{m};q^{m})_{\infty}\sum\limits_{n=0}^{\infty}\dfrac{q^{mn}}{(q;q)_{mn}(q^{mn+m};q^{m})_{\infty}}.$
Using Euler’s identity [1, p. 1919, ($2.2.5$)]
(2.2) $\dfrac{1}{(t;q)_{\infty}}=\sum\limits_{n=0}^{\infty}\dfrac{t^{n}}{(q;q)_{n}},$
(replacing $q$ by $q^{m}$ and letting $t=q^{mn+m}$) in (2.1), we get
$\sum\limits_{n=0}^{\infty}C_{m}(n)q^{n}$ $=(q^{m};q^{m})_{\infty}\sum\limits_{n=0}^{\infty}\dfrac{q^{mn}}{(q;q)_{mn}}\sum\limits_{j=0}^{\infty}\dfrac{q^{mnj+mj}}{(q^{m};q^{m})_{j}}$
$=(q^{m};q^{m})_{\infty}\sum\limits_{n,j=0}^{\infty}\dfrac{q^{mn+mnj+mj}}{(q;q)_{mn}(q^{m};q^{m})_{j}}$
$=(q^{m};q^{m})_{\infty}\sum\limits_{n,j=0}^{\infty}\dfrac{1}{m}\left(\sum\limits_{k=0}^{m-1}\zeta^{kn}_{m}\right)\dfrac{q^{n+nj+mj}}{(q;q)_{n}(q^{m};q^{m})_{j}}$
(2.3) $=\dfrac{(q^{m};q^{m})_{\infty}}{m}\sum\limits_{j=0}^{\infty}\dfrac{q^{mj}}{(q^{m};q^{m})_{j}}\sum\limits_{k=0}^{m-1}\sum\limits_{n=0}^{\infty}\dfrac{(\zeta^{k}_{m}q^{j+1})^{n}}{(q;q)_{n}}.$
Using (2.2) with $t=\zeta^{k}_{m}q^{j+1}$ in (2.3), we get
$\sum\limits_{n=0}^{\infty}C_{m}(n)q^{n}$ $=\dfrac{(q^{m};q^{m})_{\infty}}{m}\sum\limits_{j=0}^{\infty}\dfrac{q^{mj}}{(q^{m};q^{m})_{j}}\sum\limits_{k=0}^{m-1}\dfrac{1}{(\zeta^{k}_{m}q^{j+1};q)_{\infty}}$
$=\dfrac{(q^{m};q^{m})_{\infty}}{m}\sum\limits_{j=0}^{\infty}\dfrac{q^{mj}}{(q^{m};q^{m})_{j}(q^{j+1};q)_{\infty}}$
$\qquad\qquad+\dfrac{(q^{m};q^{m})_{\infty}}{m}\sum\limits_{j=0}^{\infty}\dfrac{q^{mj}}{(q^{m};q^{m})_{j}}\sum\limits_{k=1}^{m-1}\dfrac{1}{(\zeta^{k}_{m}q^{j+1};q)_{\infty}}$
$=\dfrac{1}{m}\sum\limits_{j=0}^{\infty}\dfrac{q^{mj}(q^{m(j+1)};q^{m})_{\infty}}{(q^{j+1};q)_{\infty}}$
$\qquad\qquad+\dfrac{(q^{m};q^{m})_{\infty}}{m}\sum\limits_{j=0}^{\infty}\dfrac{q^{mj}(q^{j+1};q)_{\infty}}{(q^{m};q^{m})_{j}(q^{m(j+1)};q^{m})_{\infty}}\sum\limits_{k=1}^{m-1}(\zeta^{k}_{m}q^{j+1};q)_{\infty}$
$=\dfrac{1}{m}\sum\limits_{j=0}^{\infty}q^{mj}\cdot\prod\limits_{i=m+1}^{\infty}(1+q^{i}+\ldots+q^{(m-1)i})$
$\qquad\qquad+\dfrac{1}{m}\sum\limits_{j=0}^{\infty}q^{mj}(q^{j+1};q)_{\infty}\sum\limits_{k=1}^{m-1}(\zeta^{k}_{m}q^{j+1};q)_{\infty}$
(2.4) $=\dfrac{1}{m}\sum\limits_{n=0}^{\infty}D_{m}(n)q^{n}+\dfrac{1}{m}\sum\limits_{n=0}^{\infty}E_{m}(n)q^{n},$
where the generating function of $D_{m}(n)$ clearly follows from its definition and the generating function of $E_{m}(n)$ follows from (1.7). Comparing coefficients of $q^{n}$ on both sides of (2.4) for $n>1$, we have
$C_{m}(n)=\dfrac{D_{m}(n)+E_{m}(n)}{m}$
which completes the proof of Theorem 1.4.
3. Proof of Theorem 1.5
From (1.7), it is clear that
(3.1) $\varepsilon_{3}(q)=(q;q)_{\infty}\sum\limits_{j=0}^{\infty}\dfrac{q^{3j}}{(q;q)_{j}}\left((\omega q^{j+1};q)_{\infty}+(\omega^{-1}q^{j+1};q)_{\infty}\right).$
Now, define $\chi_{3}(n)$ as follows
$\chi_{3}(n):=\omega^{n}+\omega^{-n}=\Bigg\{\begin{array}[]{lr}2\qquad\,\,\,\text{if }3\mid n,\\ -1\quad\,\,\,\,\text{if }3\nmid n.\end{array}$
Using [1, p. 1919, ($2.2.6$)] with $t=-\omega q^{j+1}$ and $t=-\omega^{-1}q^{j+1}$ in (3.1), we get
(3.2) $\varepsilon_{3}(q)=(q;q)_{\infty}\sum\limits_{j=0}^{\infty}\dfrac{q^{3j}}{(q;q)_{j}}\sum\limits_{n=0}^{\infty}\dfrac{(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}+nj}}{(q;q)_{n}}.$
Using Euler’s identity [1, p. 1919, ($2.2.5$)] with $t=q^{n+3}$ after interchanging summations in (3.2), we then have
$\varepsilon_{3}(q)$ $=(q;q)_{\infty}\sum\limits_{n=0}^{\infty}\dfrac{(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}}}{(q;q)_{n}(q^{n+3};q)_{\infty}}$
$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}}(1-q^{n+1})(1-q^{n+2})$
$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}}(1-q^{n+1}(1+q)+q^{(n+1)+(n+2)})$
$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}}-(1+q)\sum\limits_{n=0}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+2}{2}}+\sum\limits_{n=0}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+3}{2}}$
$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}}+(1+q)\sum\limits_{n=1}^{\infty}(-1)^{n}\chi_{3}(n-1)q^{\binom{n+1}{2}}$
$\qquad+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n-2)q^{\binom{n+1}{2}}$
$=(2+q)+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}}-2q(1+q)+(1+q)\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n-1)q^{\binom{n+1}{2}}$
$\qquad+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n-2)q^{\binom{n+1}{2}}$
$=2-q-2q^{2}+\sum\limits_{n=2}^{\infty}(-1)^{n}\left(\chi_{3}(n)+\chi_{3}(n-1)+\chi_{3}(n-2)\right)q^{\binom{n+1}{2}}$
$\qquad+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n-1)q^{\binom{n+1}{2}+1}.$
Now, $\chi_{3}(n)+\chi_{3}(n-1)+\chi_{3}(n-2)=0$ for all $n$. Hence,
$\varepsilon_{3}(q)=2-q-2q^{2}+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n-1)q^{\binom{n+1}{2}+1}$
which completes the proof of Theorem 1.5.
4. Proof of Theorem 1.7
Define $\chi_{m}(n)$ as follows
$\chi_{m}(n):=\sum\limits_{j=1}^{m-1}\zeta^{jn}_{m}=\Bigg\{\begin{array}[]{lr}m-1\,\,\,\text{if }m\mid n,\\ -1\quad\,\,\,\,\,\text{if }m\nmid n.\end{array}$
Using [1, p. 1919, ($2.2.6$)] with $t=-\zeta^{j}_{m}q^{n}$ in (1.7), we get
(4.1) $\varepsilon_{m}(q)$ $=(q;q)_{\infty}\sum\limits_{n=0}^{\infty}\dfrac{q^{mn}}{(q)_{n}}\sum\limits_{k=0}^{\infty}\dfrac{(-1)^{k}\chi_{m}(k)q^{\binom{k+1}{2}+nk}}{(q;q)_{k}}.$
Using Euler’s identity [1, p. 1919, ($2.2.5$)] with $t=q^{k+m}$ after interchanging summations in (4.1), we then have
$\varepsilon_{m}(q)$ $=(q;q)_{\infty}\sum\limits_{k=0}^{\infty}\dfrac{(-1)^{k}\chi_{m}(k)q^{\binom{k+1}{2}}}{(q;q)_{k}(q^{k+m};q)_{\infty}}$
(4.2) $=\sum\limits_{k=0}^{\infty}(-1)^{k}\chi_{m}(k)q^{\binom{k+1}{2}}(q^{k+1};q)_{m-1}.$
Using [1, p. $36$, ($3.3.6$)] with the substitutions N=m1N=m-1 and $z=q^{k+1}$, we get
$\varepsilon_{m}(q)$ $=\sum\limits_{k=0}^{\infty}(-1)^{k}\chi_{m}(k)q^{\binom{k+1}{2}}\sum\limits_{j=0}^{m-1}(-1)^{j}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}q^{\binom{j+1}{2}+kj}$
$=\sum\limits_{k=0}^{\infty}(-1)^{k}\chi_{m}(k)\sum\limits_{j=0}^{m-1}(-1)^{j}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}q^{\binom{j+k+1}{2}}$
$=\sum\limits_{j=0}^{m-1}(-1)^{j}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=0}^{\infty}(-1)^{k}\chi_{m}(k)q^{\binom{j+k+1}{2}}$
$=\sum\limits_{j=0}^{m-1}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=j}^{\infty}(-1)^{k}\chi_{m}(k-j)q^{\binom{k+1}{2}}$
$=-\sum\limits_{j=0}^{m-1}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=0}^{j-1}(-1)^{k}\chi_{m}(k-j)q^{\binom{k+1}{2}}$
$\qquad+\sum\limits_{j=0}^{m-1}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=0}^{\infty}(-1)^{k}\chi_{m}(k-j)q^{\binom{k+1}{2}}$
$=-\sum\limits_{j=0}^{m-1}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=0}^{j-1}(-1)^{k}\chi_{m}(k-j)q^{\binom{k+1}{2}}$
$\qquad+\sum\limits_{k=0}^{\infty}(-1)^{k}q^{\binom{k+1}{2}}\sum\limits_{j=0}^{m-1}\chi_{m}(k-j)\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}$
$=-\sum\limits_{j=0}^{m-1}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=0}^{j-1}(-1)^{k}\chi_{m}(k-j)q^{\binom{k+1}{2}}$
$\qquad+\sum\limits_{k=0}^{\infty}(-1)^{k}q^{\binom{k+1}{2}}\sum\limits_{j=0}^{m-1}\chi_{m}(k-j)\left(\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}-1\right),$
where we have used the fact that $\sum\limits_{j=0}^{m-1}\chi_{m}(k-j)=0$ in the second sum on the last line above. Hence, we have
(4.3) $\varepsilon_{m}(q)=P_{m}(q)+\sum\limits_{k=0}^{\infty}(-1)^{k}q^{\binom{k+1}{2}}\sum\limits_{j=0}^{m-1}\chi_{m}(k-j)\left(\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}-1\right),$
where $P_{m}(q):=-\sum\limits_{j=0}^{m-1}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=0}^{j-1}(-1)^{k}\chi_{m}(k-j)q^{\binom{k+1}{2}}$.
The polynomial
$\sum\limits_{j=0}^{m-1}\chi_{m}(k-j)\left(\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}-1\right)$
in (4.3) above has at most $2^{m-1}-m$ terms. Now, there are at most $\sqrt{2x}+1$ triangular numbers less than $x$. Hence, the maximum possible number of non-zero coefficients in $\varepsilon_{m}(q)$ among the first $x$ terms is at most $(2^{m-1}-m)\cdot(\sqrt{2x}+1)$ leaving the number of coefficients that vanish as less than
$x-(2^{m-1}-m)\cdot(\sqrt{2x}+1)$
which is asymptotic to $x$ as $x\rightarrow\infty$. This shows that $E_{m}(n)$ is almost always zero which completes the proof of Theorem 1.7.
5. Concluding Remarks
It would be very interesting to provide a bijective proof of Theorem 1.6.
Acknowledgments
The authors would like to thank Alexander Berkovich and Rahul Kumar for their kind interest.